Say true or false and justify your answer:   3^0 = (1000)^0

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us consider LHS = 3^0
=1
Now, consider RHS = 1000^0
=1
By comparing LHS and RHS
LHS = RHS
3^0= 1000^0
Hence, the given statement is true

Thanks